# A Refreshed Similarity-based Upsampler for Direct High-Ratio Feature Upsampling

### v1

@@ 4.1 Explicitly Controllable Query-Key Alignment
From [54, 40], an HR guidance feature $\bm{y}$ from a shallow layer usually captures more detail-related information while the LR deep feature $\bm{x}$ encodes more semantic-related information. Thus, we informally refer to $\bm{y}$ (and its linear projection, the query feature $\bm{q}$) and $\bm{x}$ (and its linear projection, the key feature $\bm{k}$) as residing in the detail space and the semantic space, respectively, as presented in Fig. 3.
@@ 4.1.1 Semantic-Aware Mutual-Alignment
Remark 1: From the visualization in Fig. 2 (b), we can clearly observe that: 1) compared to the original linearly-projected query $\bm{q}$, the explicitly optimized query $\bm{q}_{GF}$ exhibits a stronger semantic resemblance to the key $\tilde{\bm{k}}$. This can promote more accurate semantic-aware similarity computation of $\bm{s}_{s}$ for better performance (see Sec. 5.3.2); 2) $\bm{q}_{GF}$ effectively preserves the original intricate structures in $\bm{q}$, which would guide the upsampling to achieve a higher structural fidelity in the upsampled feature $\tilde{\bm{x}}$ (also see Fig. 16). All these advantages finely comply with our design motivation for Eq. (4) and substantiate its rationality.
@@ 4.1.2 Detail-Aware Self-Alignment
Remark 2: 1) The rationality of the detail-aware self-alignment can be explained from another perspective. Specifically, it is well-acknowledged that within a small local region of a relatively shallow feature, if two pixels exhibit similarity in detail structures, there is a high likelihood that they should possess similar semantics in the upsampled feature $\tilde{\bm{x}}$ [58]. Thus, it is reasonable to exploit the self-similarity on query $\bm{q}$ itself for guiding the upsampling process; 2) The incorporation of the detail-aware similarity favorably encourages fine-grained detail preservation and the smoothed key generation design would indeed bring some performance improvement as validated in Sec. 5.3.2.
@@ 4.2.1 Paired Central Difference Convolution
For a better understanding of the working mechanism of the proposed PCDC, we provide a visual illustration of Eq. (11) with the kernel size $K$ as 3. As observed from Fig. 4, through the learnable convolution weights, the differences between the central query element $\bar{\bm{q}}_{i}$ and its neighboring key elements $\bar{\bm{k}}_{j_{n}}$ can be well captured and flexibly processed. Naturally, the PCDC output $\bm{v}$ has the capability to encode the local similarity information between $\bar{\bm{q}}$ and $\bar{\bm{k}}$.
@@
We can find that the proposed PCDC inherently consists of a vanilla grouped convolution on $\bar{\bm{k}}$ with weight $\bm{w}$ (the first term in Eq. (12)), and a $1\times 1$ convolution on $\bar{\bm{q}}$ with the weight as the aggregation of $\bm{w}$ in the spatial dimension (the second term in Eq. (12)). These operations can be easily and efficiently implemented based on PyTorch [59] (see the pseudocode in the supplementary material).
@@ 4.2.2 PCDC-Block for Similarity Calculation
Figure 7: Illustration of the query-key similarity calculation process under different neighborhood selection strategies for $\mathcal{N}\left(i\right)$: (a) without FNS and (b) with FNS. Here the kernel size $K=1$ and the upsampling ratio is 2. Pixels with the same color and shape in the query and key are paired for computing the similarity scores.
@@ 4.3 Fine-grained Neighbor Selection
To understand our FNS strategy more intuitively, Fig. 7 presents the computation process of similarity score for the query-key pair under a simplified case where the kernel size $K=1$ and each query pixel would pair with only one neighboring pixel in the key feature. As seen, for the case without FNS, all the four HR query elements in the red dotted box would pair with the same neighboring LR key element to compute the similarity score. However, for our proposed FNS, the neighbor selection is executed on the bilinearly upsampled HR key feature rather than the original LR one. In this manner, the four HR query elements would separately pair with different neighboring HR key elements, as marked by four red dotted boxes. Such a smoother query-key pair selection manner would promote the smoothness of similarity scores for adjacent HR query pixels with different neighbors in the key feature. These advantageous characteristics of FNS can be readily extended to scenarios with larger kernel sizes. Albeit simple, this FNS strategy can be implemented efficiently without introducing any additional computational cost, and can fundamentally eliminate mosaic artifacts, which can be observed by comparing the upsampled features $\tilde{\bm{x}}$ in Figs. 2 (a) and (b). Please refer to Fig. 17 for more experimental validation.
@@
Remark 4: Overall, for the three key components in Eqs. (2)(3), we systematically reform the similarity-based upsampling methodology. Specifically, we explicitly optimize the query-key alignment for facilitating the computation of $\bm{s}_{i}$, design a flexible similarity function $\text{sim}(\cdot,\cdot)$, and develop a fine-grained neighbor selection strategy for $\mathcal{N}(i)$ as derived in Secs. 4.1, 4.2, and 4.3, respectively. We accordingly construct a refreshed similarity-based feature upsampler, named ReSFU (see Fig. 2 (b)). Notably, the proposed three designs harmoniously form an inseparable whole with each part being indispensable (see Sec. 5.3.2). The specific and meticulous design of each part ultimately enables ReSFU to be applicable to different network structures, which will be validated in the experiments below.
@@ 5 Experiments
In this section, we comprehensively substantiate the applicability of our proposed ReSFU by conducting extensive experiments on various network structures for the semantic segmentation task. Then, we provide detailed model verification and ablation studies to clearly show the working mechanism of ReSFU and validate the rationality of every involved component. Finally, in order to better demonstrate the universality of the proposed ReSFU, we extend it to more application scenarios with a variety of datasets, including medical image segmentation, instance segmentation and panoptic segmentation.
@@ 5.1 Experimental Setup
Implementation details. For similarity calculation in Eq. (12), the number of output channels $L$ and groups $G$ are 32 and 4, respectively. The projection dimension $D$ is 32 and the kernel size $K$ is 3. For all the involved networks, we modify the default upsampling method (nearest neighbor for FPN neck [60] or bilinear for other networks) by substituting it with alternative upsamplers (replacement details are described in the experimental comparisons below), while keeping other training settings constant for end-to-end training on NVIDIA V100 GPUs. Please refer to the supplementary material for more details.
@@
Figure 9: For the CNN-based PSPNet with non-hierarchical upsampling, (a) previous baselines adopt a two-step upsampling by default; (b) Our ReSFU performs a one-step ×$\times 4$ upsampling. The guidance feature is from the involved stem ConvBlock layer in the ResNet backbone.
@@ 5.2 Semantic Segmentation on Various Architectures
In this section, for the semantic segmentation task, we verify the effectiveness of our proposed ReSFU by conducting comprehensive experiments on various network structures with different types of popular backbones including ResNet [61], Vision TransFormer (ViT) [40], Mix Transformer (MiT) [1], and Swin Transformer [62]. For the selection of network structures, we consider not only the hierarchical upsampling manner adopted by the existing baselines, e.g., SegFormer [1], but also the non-hierarchical upsampling manner, e.g., PSPNet [37] and Segmenter [3], as well as MaskFormer with the FPN head [51], covering most of the existing architecture types involving feature upsampling. Following [2, 13], the semantic segmentation experiments are evaluated using the single-scale testing on the ADE20K dataset [63], which is a widely adopted large-scale scene parsing dataset containing over 20,000 images with pixel-level annotations for 150 object categories. Three typical metrics are used for quantitative evaluation, including mean IoU (mIoU), mean pixel accuracy (mAcc), and boundary IoU (bIoU) [64]. For the bIoU computation, the pixel distance is set to $2\%$ [64]. More experiments on more datasets are provided in the supplementary material.
@@
5.2.1 Evaluation on PSPNet with ResNet Backbone
Upsampling Details. We first select the representative PSPNet [37] with the pioneering CNN-based backbone ResNet [61] for the semantic segmentation task. Fig. 9 presents the replacement manners of different upsampling methods. As seen, for the feature $\bm{x}$ extracted in the last block, to achieve the $\times$4 upsampling for the final segmentation output, previous upsamplers adopt two successive $\times$2 upsampling processes by default. In contrast, our ReSFU is simply embedded with a direct $\times$4 upsampling. For guidance-based upsamplers, e.g., FADE, SAPA, and ReSFU, the guidance feature is from the stem convolution block.
Figure 10: For the Transformer-based Segmenter with non-hierarchical upsampling, (a) previous methods adopt two ×$\times 2$ upsampling processes; (b) Our ReSFU directly performs a ×$\times 4$ upsampling. The guidance HR feature is obtained by applying an additional ConvBlock to the input image.
Performance Comparison. Table I reports the quantitative results on ADE20K with two variants of the ResNet backbone, i.e., ResNet50 and ResNet101. It is clearly observed that with the comparable number of network parameters, our proposed ReSFU consistently outperforms other upsampling approaches in three metrics, i.e., mIoU/mAcc/bIoU (%). Taking PSPNet-ResNet101 as an example, Fig. 8 presents the upsampled features $\widetilde{\bm{x}}$ and the corresponding segmentation results obtained with different upsamplers. [^2] For feature visualization in all the experiments, we follow [22] and adopt principal component analysis [65] to reduce the features to three channels corresponding to the RGB values. Note that since the feature extractors are not frozen and are trained together with different upsamplers, there may be significant variations in color tone across different feature visualizations. From the feature visualizations, we can find that the object boundary upsampled by our ReSFU is much clearer, thus boosting more accurate segmentation results. In comparison with other approaches with multi-step ×$\times 2$ upsampling processes, even with a direct one-step high-ratio upsampling, our proposed ReSFU can still obtain superior performance, which facilitates the deployment in practical applications.
@@ 5.2.1 Evaluation on PSPNet with ResNet Backbone
5.2.2 Evaluation on Segmenter with ViT Backbone
Upsampling Details. Here we introduce the representative ViT-based network, Segmenter [3] with the FCN-head [4], and present the incorporation manners of different feature upsampling methods as depicted in Fig. 10. Similar to PSPNet, other baselines are inserted in an iterative ×$\times 2$ upsampling manner while the proposed ReSFU is operated with a direct high-ratio upsampling. To generate the guidance, we apply an extra ConvBlock to the input image in order to obtain a shallow HR feature. The ConvBlock consists of “Conv3$\times$3+GroupNorm+ReLU+Conv3$\times$3” layers, with the channel dimension being “$3\to 64\to 32$”.
Figure 12: For SegFormer with hierarchical upsampling, (a) previous baselines execute the iterative ×$\times 2$ upsampling with utilizing $C3$, $C2$, and $C1$ as the guidance feature, respectively; (b) Our ReSFU only adopts the guidance $C1$ for the direct upsampling with different ratios.
@@ 5.2.2 Evaluation on Segmenter with ViT Backbone
Performance Comparison. For comprehensiveness, we evaluate our ReSFU based on the ViT backbone with different sizes, e.g., ViT-S and ViT-L. As reported in Table II, the proposed ReSFU evidently surpasses other methods, and achieves the best mIoU/mAcc/bIoU (%) scores under different backbones with the comparable number of parameters. The visual comparison is presented in Fig. 11, clearly showing that our proposed ReSFU achieves better feature upsampling effects with more accurate delineation of object boundaries, e.g., the desk lamp. These findings effectively validate the suitability of our ReSFU in the current popular Transformer-based architectures, demonstrating its strong potential for practical application.
5.2.3 Evaluation on SegFormer with MiT
Upsampling Details. For the semantic segmentation task, SegFormer [1] is another mainstream Transformer-based network structure. Different from Segmenter with a non-hierarchical upsampling architecture, the upsampling process in SegFormer is composed of hierarchical forms among different levels of features. As displayed in Fig. 12, the encoder outputs four levels of features, i.e., $C1$, $C2$, $C3$, and $C4$ with different resolutions. The LR features $C2$, $C3$, and $C4$ require a ×$\times 2$, ×$\times 4$, and ×$\times 8$ upsampling, respectively, to be concatenated with the HR $C1$ together for the final segmentation. For the existing upsamplers, they iteratively perform the multi-step ×$\times 2$ upsampling. For instance, to upsample the LR feature $C4$ to the same resolution as $C1$, the guidance-based upsampling methods, e.g., FADE and SAPA, sequentially execute three ×$\times 2$ upsampling procedures by adopting the adjacent-level features $C3$, $C2$ and $C1$ as guidance, respectively. However, for our ReSFU, no matter which level the feature is at, only one direct high-ratio upsampling is utilized under the fixed guidance of $C1$, thus halving the total number of upsampling modules.
Figure 14: Visual comparison of the upsampled features $P2$ in Fig. 15 and predicted masks with MaskFormer-Swin-L.
Figure 15: For MaskFormer with the FPN neck, (a) previous methods sequentially take $C3$, $C2$, and $C1$ as guidance to execute the ×$\times 2$ upsampling; (b) Our ReSFU directly utilizes $C1$ to guide the upsampling for different levels of features.
Performance Comparison. Table III lists the quantitative results on two variants of the backbone MiT, i.e., MiT-B1 and MiT-B5, and the proposed ReSFU achieves better overall performance on all these metrics. Fig. 13 illustrates the ×$\times 8$ upsampled feature $\widetilde{C4}$ and the segmentation result obtained by different upsamplers. We can observe that for the comparing approaches with three consecutive ×$\times 2$ upsampling modules, e.g., CARAFE and DySample, the final upsampled features exhibit blurry effects. Nevertheless, even with a simpler ×$\times 8$ upsampling module, our ReSFU still accomplishes better detail preservation without mosaic artifacts. These favorable effects are mainly attributed to the careful designs of the controllable feature alignment and the fine-grained neighbor selection process, which enables ReSFU to be suitable for this direct high-ratio upsampling scenario. Thus, different from the current guidance-based feature upsampling methods, our ReSFU does not require features of adjacent levels as guidance, which naturally makes it possible to be extended to more network architectures not limited to hierarchical upsampling.
5.2.4 Evaluation on MaskFormer with Swin-Transformer
Upsampling Details. MaskFormer [51] is also a cutting-edge Transformer-based network for the segmentation task. It contains a FPN neck-based upsampling module [60], different from that adopted in Segmenter and SegFormer. As presented in Fig. 15, the FPN neck consists of three ×$\times 2$ upsampling procedures. For the guidance-based comparing methods, they regard the adjacent-level features, i.e., $C3$, $C2$, and $C1$, as the guidance for these three ×$\times 2$ upsampling modules, respectively. However, for ReSFU, we always utilize the relatively shallower feature $C1$ as guidance to introduce more information with details.
@@ 5.2.4 Evaluation on MaskFormer with Swin-Transformer
Performance Comparison. First, we quantitatively evaluate all the comparing methods based on two versions of Swin Transformer [62] with varying sizes, including Swin-B and Swin-L. As observed from Table IV, ReSFU consistently obtains state-of-the-art outcomes over all metrics. Then based on MaskFormer-Swin-L, we visualize the HR features upon upsampling $P2$. From the upper row in Fig. 14, it is clearly seen that the comparing methods suffer from either blurriness or checkerboard artifacts [66]. In contrast, our ReSFU can generate clearer object boundaries and preserve more useful content, which in turn promotes better mask prediction as shown in the lower row.

### v2

@@ 4.1 Explicitly Controllable Query-Key Alignment
From [26, 27], the HR guidance feature $\bm{y}$ from a shallow layer usually captures more detail-related information while the LR deep feature $\bm{x}$ encodes more semantic-related information. Thus, we informally refer to $\bm{y}$ (and its linear projection, the query feature $\bm{q}$) and $\bm{x}$ (and its linear projection, the key feature $\bm{k}$) as residing in the detail space and the semantic space, respectively, as presented in Fig. 3.
@@ 4.1.1 Semantic-Aware Mutual-Alignment
Remark 1: From the visualization in Fig. 2 (b), we can clearly observe that: 1) compared to the original linearly-projected query $\bm{q}$, the explicitly optimized query $\bm{q}_{GF}$ exhibits a stronger semantic resemblance to the key $\tilde{\bm{k}}$. This can promote more accurate semantic-aware similarity computation of $\bm{s}_{s}$ for better performance (see Sec. 5.3.2); 2) $\bm{q}_{GF}$ effectively preserves the original intricate structures in $\bm{q}$, which would guide the upsampling to achieve a higher structural fidelity in the upsampled feature $\tilde{\bm{x}}$ (also see Fig. 13). All these advantages finely comply with our design motivation for Eq. (4) and substantiate its rationality. Please note that although guided filter is the existing technique, we are the first to propose the concept of semantic-aware mutual alignment based on the essential understanding of the pipeline as analyzed in Sec. 3. Driven by the novel concept, the guided filter is indeed the most natural solution for efficient feature alignment in semantic space.
@@ 4.1.2 Detail-Aware Self-Alignment
Remark 2: 1) The rationality of the detail-aware self-alignment can be explained from another perspective. Specifically, it is well-acknowledged that within a small local region of a relatively shallow feature, if two pixels exhibit similarity in detail structures, there is a high likelihood that they should possess similar semantics in the upsampled feature $\tilde{\bm{x}}$ [32]. Thus, it is reasonable to exploit the self-similarity on query $\bm{q}$ itself for guiding the upsampling process; 2) The incorporation of the detail-aware similarity favorably encourages fine-grained detail preservation and the smoothed key generation design would indeed bring some performance improvement as validated in Sec. 5.3.2. It is worth mentioning that this proposed novel self-alignment technique can be easily transferred to the existing upsampling methods for better performance, as verified in SM.
@@ 4.2.1 Paired Central Difference Convolution
The closer $\bar{\bm{k}}_{j_{n}d}-\bar{\bm{q}}_{id}$ is to 0, the more similar $\bar{\bm{q}}_{id}$ and $\bar{\bm{k}}_{j_{n}d}$ are. From [33], this difference form focuses on capturing center-oriented gradient information which carries semantic information and can contribute to generating meaningful upsampling kernels for boosting the higher-quality generation of HR features. For a better understanding of the working mechanism of the proposed PCDC, we provide a visual illustration of Eq. (11) with the kernel size $K$ as 3. As observed from Fig. 4, through the learnable convolution weights, the differences between the central query element $\bar{\bm{q}}_{i}$ and its neighboring key elements $\bar{\bm{k}}_{j_{n}}$ can be well captured and flexibly processed. Naturally, the PCDC output $\bm{v}$ has the capability to encode the local similarity information between $\bar{\bm{q}}$ and $\bar{\bm{k}}$.
@@
We can find that the proposed PCDC inherently consists of a vanilla grouped convolution on $\bar{\bm{k}}$ with weight $\bm{w}$ (the first term in Eq. (12)), and a $1\times 1$ convolution on $\bar{\bm{q}}$ with the weight as the aggregation of $\bm{w}$ in the spatial dimension (the second term in Eq. (12)). These operations can be easily and efficiently implemented based on PyTorch [34] (see the pseudocode in SM).
@@ 4.2.2 PCDC-Block for Similarity Calculation
Then the complete calculation procedures for semantic-aware similarity scores in Eq. (7) and detail-aware similarity scores in Eq. (8) are:
[…]
where the operator PCDC-Block is illustrated by Fig. 5; $\bm{\theta}_{s}$ and $\bm{\theta}_{d}$ represent the network parameters involved in GroupNorm, PCDC, and the channel compressor for computing $\bm{s}_{s}$ and $\bm{s}_{d}$, respectively.
@@
Figure 7: Illustration of the query-key similarity calculation process under different neighbor selection strategies for $\mathcal{N}(i)$: (a) without FNS and (b) with FNS. Here the kernel size $K=3$, and the neighbors $\mathcal{N}(i)$ of the highlighted red pixel $i$ are $j_{1},\cdots,j_{9}$. Pixels marked in circles with the same color in the query and key features are paired for computing the similarity scores.
@@ 4.3 Fine-grained Neighbor Selection
To understand our FNS strategy more intuitively, Fig. 7 presents the query-key similarity calculation process with the kernel size $K=3$ for a ×$\times 2$ upsampling. As seen, for the case without FNS, the four HR query elements with different colors in a $2\times 2$ local region share the same set of neighboring LR key pixels indexed by $j_{1},\cdots,j_{9}$. However, for our proposed FNS, the neighbor selection is executed on the HR key feature. In this manner, the four HR query elements in a $2\times 2$ local region would separately pair with different nine neighboring HR key elements. Such a smoother query-key pair selection manner would promote the smoothness of similarity scores for adjacent HR query pixels with different neighbors in the key feature. Albeit simple, this FNS strategy can be implemented efficiently without introducing any additional computational cost, and can fundamentally eliminate mosaic artifacts, which can be observed by comparing the upsampled features $\tilde{\bm{x}}$ in Figs. 2 (a) and (b). Please refer to Fig. 14 for more experimental validation.
@@
Algorithm 1 Algorithm of ReSFU for feature upsampling
1:LR deep feature $\bm{x}\in\mathbb{R}^{hw\times C}$, HR guidance feature $\bm{y}\in\mathbb{R}^{HW\times c}$; parameters of linear projections $\bm{\theta}_{\bm{k}}$ and $\bm{\theta}_{\bm{q}}$, hyperparamters in guided filter $r$ and $\epsilon$, parameters of the PCDC-Blocks $\bm{\theta}_{s}$ and $\bm{\theta}_{d}$.
2:Upsampled deep feature $\tilde{\bm{x}}\in\mathbb{R}^{HW\times C}$.
3:$\bm{k},\bm{q}\leftarrow\text{Proj}_{\bm{\theta}_{\bm{k}}}(\bm{x}),\text{Proj}_{\bm{\theta}_{\bm{q}}}(\bm{y})$triangleright linear projection
[…]
9:$\bm{s}\leftarrow\bm{s}_{s}+\bm{s}_{d}$triangleright Eq. (9)
10:$\tilde{\bm{x}}_{i}\leftarrow\text{Softmax}(\bm{s}_{i})\bm{x}_{\mathcal{N}(i)};\mathcal{N}(i)$: w/ FNS $\triangleright$ Eq. (2), Fig. 7 (b)
4.4 Overall Framework of ReSFU
Overall, for the three key components in Eqs. (2)(3), we systematically reform the similarity-based upsampling methodology. Specifically, we explicitly optimize the query-key alignment for facilitating the computation of $\bm{s}_{i}$, design a flexible similarity function $\text{sim}(\cdot,\cdot)$, and develop a fine-grained neighbor selection strategy for $\mathcal{N}(i)$ as derived in Secs. 4.1, 4.2, and 4.3, respectively. We accordingly construct a refreshed similarity-based feature upsampler, named ReSFU. The overall framework of ReSFU is illustrated in Fig. 2 (b) and the complete computation procedure is summarized in Alg. 1. Notably, the proposed three designs harmoniously form an inseparable whole with each part being indispensable (see Sec. 5.3.2). The specific and meticulous design of each part ultimately enables ReSFU to be applicable to different network structures, which will be validated in Sec. 5 below.
Remark 4: Please be aware that we only continue the similarity-based research paradigm introduced by SAPA. Nevertheless, we have made highly comprehensive and systematic fundamental enhancements to SAPA across multiple aspects, including problem formulation, essence discovery, and methodological designs. All these efforts promote our proposed ReSFU to possess wider application scenarios than SAPA. These merits will be fully validated through extensive experiments based on diverse configurations.
@@ 5 Experiments
In this section, we comprehensively substantiate the applicability of our proposed ReSFU by conducting extensive experiments on various network structures for the semantic segmentation task. Then, we provide detailed model verification and ablation studies to clearly show the working mechanism of ReSFU and validate the rationality of every involved component. Finally, in order to better demonstrate the universality of the proposed ReSFU, we extend it to more application scenarios with a variety of datasets, including medical image segmentation, instance segmentation, panoptic segmentation, object detection, and monocular depth estimation.
@@ 5.1 Experimental Setup
Implementation details. For similarity calculation in Eq. (12), the number of output channels $L$ and groups $G$ are 32 and 4, respectively. The projection dimension $D$ is 32 and the kernel size $K$ is 3. For all the involved networks, we modify the default bilinear/nearest neighbor upsampling method by substituting it with alternative upsamplers (replacement details are described in the experimental comparisons below), while keeping other training settings constant for end-to-end training on NVIDIA V100 GPUs. Please refer to SM for more details.
Figure 8: Illustration of different backbones, i.e., (a) PSPNet, (b) Segmenter, (c) SegFormer, and (d) MaskFormer (FPN) with iterative (upper row) and direct (lower row) upsampling manners. Please refer to SM for more details.
@@ 5.2 Semantic Segmentation on Various Architectures
5.2.1 Experimental Settings
In this section, for the semantic segmentation task, we verify the effectiveness of our proposed ReSFU by conducting comprehensive experiments on various network structures with different types of popular backbones including ResNet [35], Vision TransFormer (ViT) [27], Mix Transformer (MiT) [1], and Swin Transformer [36]. For the selection of network structures, we consider not only the hierarchical upsampling manner adopted by the existing baselines, e.g., SegFormer [1], but also the non-hierarchical upsampling manner, e.g., PSPNet [37] and Segmenter [3], as well as MaskFormer with the FPN head [38], covering most of the existing architecture types involving feature upsampling. Fig. 8 illustrates the corresponding network structures of different backbones, which contain two different upsampling manners, i.e., iterative one and direct one. Please note that for other baselines, we implement them in an iterative upsampling manner by default for fair comparison. For our ReSFU, we specifically implement it in the more difficult direct upsampling manner to validate its capability in achieving the direct high-ratio upsampling. Actually, directly inserting other comparison methods in the direct upsampling manner would impair their performance. Please refer to Sec. 5.5 and SM for more discussion and implementation details.
Following [2, 16], the semantic segmentation experiments are evaluated using the single-scale testing on the ADE20K dataset [39], which is a widely adopted large-scale scene parsing dataset containing over 20,000 images with pixel-level annotations for 150 object categories. Three typical metrics are used for quantitative evaluation, including mean IoU (mIoU), mean pixel accuracy (mAcc), and boundary IoU (bIoU) [40]. For the bIoU computation, the pixel distance is set to $2\%$ [40]. Experiments on more datasets are provided in SM.
@@ 5.2.1 Experimental Settings
5.2.2 Experimental Results
Evaluation on PSPNet with ResNet backbone. We first select the representative PSPNet [37] with the pioneering CNN-based backbone ResNet [35] for the semantic segmentation task. Fig. 8 (a) presents the iterative and direct upsampling manners for PSPNet.
@@ 5.2.2 Experimental Results
Table I reports the quantitative results on ADE20K with two variants of the ResNet backbone, i.e., ResNet50 and ResNet101. It is clearly observed that with the comparable number of network parameters, our proposed ReSFU consistently outperforms other upsampling approaches in three metrics, i.e., mIoU/mAcc/bIoU (%). Taking PSPNet-ResNet101 as an example, Fig. 9 presents the upsampled features $\widetilde{\bm{x}}$ and the corresponding segmentation results obtained with different upsamplers. [^2] For feature visualization in all the experiments, we follow [24] and adopt principal component analysis [41] to reduce the features to three channels corresponding to the RGB values. Note that since the feature extractors are not frozen and are trained together with different upsamplers, there may be significant variations in color tone across different feature visualizations. From the feature visualizations, we can find that the object boundary upsampled by our ReSFU is much clearer, thus boosting more accurate segmentation results. In comparison with other approaches with multi-step ×$\times 2$ upsampling processes, even with a direct one-step high-ratio upsampling, our proposed ReSFU can still obtain superior performance, which facilitates the deployment in practical applications.
@@
Evaluation on Segmenter with ViT backbone. Here we introduce the representative ViT-based network, Segmenter [3] with the FCN-head [4], and present the iterative/direct upsampling manners as depicted in Fig. 8 (b).
For comprehensiveness, we evaluate our ReSFU based on the ViT backbone with different sizes, e.g., ViT-S and ViT-L. As reported in Table II, the proposed ReSFU evidently surpasses other methods, and achieves the best mIoU/mAcc/bIoU (%) scores under different backbones with comparable number of parameters. The visual comparison is presented in Fig. 10, clearly showing that ReSFU achieves better feature upsampling effects with more accurate delineation of object boundaries, e.g., the desk lamp. These findings effectively validate the suitability of our ReSFU in the current popular Transformer-based architectures, demonstrating its strong potential for practical application.
@@
Evaluation on SegFormer with MiT. For the semantic segmentation task, SegFormer [1] is another mainstream Transformer-based network structure. Different from Segmenter with a non-hierarchical upsampling architecture, the upsampling process in SegFormer is composed of hierarchical forms among different levels of features, as displayed in Fig. 8 (c).
@@
Figure 12: Visual comparison of the upsampled features $P2$ in Fig. 8 (d) and predicted masks with MaskFormer-Swin-L.
Table III lists the quantitative results on two variants of the backbone MiT, i.e., MiT-B1 and MiT-B5, and the proposed ReSFU achieves better overall performance on all these metrics. Fig. 11 illustrates the ×$\times 8$ upsampled feature $\widetilde{C4}$ and the segmentation result obtained by different upsamplers. We can observe that for the comparing approaches with three consecutive ×$\times 2$ upsampling modules, e.g., CARAFE and DySample, the final upsampled features exhibit blurry effects. Nevertheless, even with a simpler ×$\times 8$ upsampling module, our ReSFU still accomplishes better detail preservation without mosaic artifacts. These favorable effects are mainly attributed to the careful designs of the controllable feature alignment and the fine-grained neighbor selection process, which enables ReSFU to be suitable for this direct high-ratio upsampling scenario. Thus, different from the current guidance-based feature upsampling methods, our ReSFU does not require features of adjacent levels as guidance, which naturally makes it possible to be extended to more network architectures not limited to hierarchical upsampling.
@@
Evaluation on MaskFormer with Swin-Transformer. MaskFormer [38] is also a cutting-edge Transformer-based network for the segmentation task. As presented in Fig. 8 (d), it contains a FPN neck-based upsampling module [42], different from that adopted in Segmenter and SegFormer.
First, we quantitatively evaluate all the comparing methods based on two versions of Swin Transformer [36] with varying sizes, including Swin-B and Swin-L. As observed from Table IV, ReSFU consistently obtains state-of-the-art outcomes over all metrics. Then based on MaskFormer-Swin-L, we visualize the HR features upon upsampling $P2$. From the upper row in Fig. 12, it is clearly seen that the comparing methods suffer from either blurriness or checkerboard artifacts [43]. In contrast, our ReSFU can generate clearer object boundaries and preserve more useful content, which in turn promotes better mask prediction as shown in the lower row.